OpenStudy (anonymous):

i dont get this one? Choose the correct simplification of x^8 y^13\x^3 y^8

OpenStudy (anonymous):

\[\frac{x^{8}y^{13}}{x^{3}y^{8}}\] is it this?

OpenStudy (anonymous):

Yesss

OpenStudy (apoorvk):

\[\frac {a^m}{a^n} = a^{m-n}\] do that for both 'x' and 'y'

OpenStudy (anonymous):

law of indices states when dividing indices u hve to subtract.

OpenStudy (anonymous):